In a survey of 458 random households, people take pride in discussing their methods of environmental protection. 236 manage their trash by separating biodegradables from non-biodegradable. Biodegradables are converted into fertilizers while non-biodegradable are disposed of properly. Use a 95% confidence to estimate the proportion of all households who have good practices of trash management.

p = 

E = 

Lower Confidence Limit = 

Upper Confidence Limit = 


1
Expert's answer
2022-03-08T08:19:07-0500

"p=\frac{236}{458}=0.5153."

"E=\sqrt{\frac{0.5152(1-0.5153)}{458}}=0.0234."

Lower Confidence Limit"=0.5153-1.96*0.0234=0.4694."

Upper Confidence Limit"=0.5153+1.96*0.0234=0.5612."